OpenStudy (anonymous):

Find the solution, given the replacement set. 5x - 2y = 6 {(-2, -8), (0, -3), (1, -1), (2, 4), (4, 7)}

hero (hero):

Well, let's see... 5(-2)-2(-8) = -10+16 = 6 5(0)-2(-3) = 6 5(1)-2(-1) = 5+2 = 7 5(2)-2(4) = 10-8 = 2 5(4)-2(7) = 20-14 = 6 so take your pick

hero (hero):

The first two sets of coordinates as well as the last one produce a true solution, so...

hero (hero):

Well, what is the explanation for this?